OpenStudy (anonymous):

Describe the relationship between the number 2 x 104 and 4 x 106. A) 4 x 106 is twice 2 x 104 B) 4 x 106 is 20 times more than 2 x 104 C) 4 x 106 is 200 times more than 2 x 104 D) 4 x 106 is 2000 times more than 2 x 104

Directrix (directrix):

Are these supposed to be: 2 x 10^4 and 4 x 10^6 ?

Directrix (directrix):

@Doll.face

OpenStudy (jiteshmeghwal9):

\[2 \times 10^4\]\[4 \times 10^6\]

OpenStudy (jiteshmeghwal9):

@Doll.face is this ur question ?

Directrix (directrix):

If so, 2 x 10^4 4 x 10^6 = 2 * 2 * 10^2 * 10^4 4 x 10^6 = 2 * 10^4 * ( 2 * 10^2 ) 4 * 10 ^6 is ( 2 * 10^2 ) times more than 2 * 10^4

Directrix (directrix):

How much is this equal: ( 2 * 10^2 ) @Doll.face

OpenStudy (anonymous):

( 4*10^6 )/( 2*10^4 ) = 200 C.

OpenStudy (anonymous):

I love how they all just solved it among themselves